lead has a law specific heat capacity and is easily melted. How many joules of energy are needed to raise the temp of 2* 10^3 of Pb(s) from 298k to its melting point?
2) A 1.0lb of gold ingot must be heated from 29c to its melting point. Calculate the energy required.
1
Expert's answer
2014-01-29T12:32:41-0500
The melting point of Pb is 600,61 K.
Q=cmdT
Mass of Pb:
m = 2*10^3 Kg,
Temperature difference Δ T = 600,61К - 298К = 302,61К
Specific Heat of Pb = 0.13 кJ/kg*K
Q=0.13*2*10^3*302.61=78678.6 kJ

2) The melting point of Au is 1337.33 K.
Q=cmdT
1 lb = 453.6 g
Mass of Au m = 0.4536 kg,
Temperature difference Δ T = 1337.33 К - (29+273) = 1035.33 К
Specific Heat of Au = 0.129 кJ/kg*K
Q=0.129*0.4536*1035.33=60.58 kJ